OpenStudy (anonymous):

An equation is shown below: 8(2x – 14) + 13 = 4x – 27 Part A: Write the steps you will use to solve the equation, and explain each step. Part B: What value of x makes the equation true?

OpenStudy (hockeychick23):

8(2x – 14) + 13 = 4x – 27 First distribute the 8 on the left hand side: 8*2x=16x -14*8= -112 So now on the left hand side we have: 16x-112+13 Now we simplify that by adding like terms: 16x-99 So now we have: 16x-99=4x-27 Next we subtract 4x from both sides: 16x-4x-99=4x-4x-27 So we have 12x-99=-27 so we need to add 99 to both sides to get x alone. 23x-99+99= -27+99 23x=72 Divide both sides by 23 X= 3 3/23 or 72/23

OpenStudy (anonymous):

@hockeychick23 Can you help me with part b

OpenStudy (anonymous):

@Compassionate Can you help me with part B ?

OpenStudy (hockeychick23):

X= 3 3/23 or 72/23 is part B

OpenStudy (anonymous):

Thank you

OpenStudy (hockeychick23):

no problem :)